A thin layer of air is between 2 pieces of glass.
Answer #1

t = thickness

\lambdagreen = 520

\lambdablue = 433.33

for reflected back strongly ,:

2t = n \lambdagreen                               eq-1

for darkness :

2t = (n + 0.5) \lambdablue                               eq-2

using eq-1 and eq-2

n \lambdagreen = (n + 0.5) \lambdablue

n (520) = (n + 0.5) (433.33)

n = 2

using eq-1

2t = n \lambdagreen   

2 t = 2 (520)

t = 520 nm